OpenStudy (anonymous):

Find a particular solution to: y'' + 6y' + 84 = -7texp(5t)

OpenStudy (anonymous):

*Correction: y'' + 6y' + 8y = -7texp(5t)

OpenStudy (anonymous):

the A.E. IS D^2 +6D +8=0 => D= -2,-4 Yc= c1e^-2x +c2e^-4x

OpenStudy (anonymous):

paricular integral (p.I.) =1/ (D^2 +6D +8) * -7t *e(5t) do u know how to go from here???

OpenStudy (anonymous):

Yes, I got that part. I'm stuck on the particular solution part because webwork says my answer is wrong. Thanks for your thelp though. I appreciate it.